.

If you want to fit a linear regression line y=a+bx from the above data, then what is the value of 'a'?

Single choice.


.

If you want to fit a linear regression line y=a+bx from the above data, then what is the value of 'a'?

Single choice.






1
Expert's answer
2021-07-26T14:35:31-0400

A linear regression line has an equation of the form: y = a+ bx

where 'Y' is called the dependent variable or response.

and 'X' is called independent variable or predictors or explanatory variable

'a' is the y-intercept of the line and

'b' is the slope of the line.




From the above data,

"\sum x = 247\\\sum y=486\\\sum xy=20485\\\sum x^2=11409\\\sum y^2=40022\\n=6"


Based on the above table, the following is calculated:

"\bar X=\dfrac{1}{n}\sum X_i=\dfrac{247}{6}=41.167"


"\bar Y= \dfrac{1}{n}\sum Y_i=\dfrac{486}{6}=81\\\ \\SS_{XX}=\sum X_i^2-\dfrac{1}{n}(\sum X_i)^2=11409-\dfrac{(247)^2}{6}=120.83"


"SS_{YY}=\sum Y_i^2-\dfrac{1}{n}(\sum Y_i)^2=40022-\dfrac{(486)^2}{6}=656"


"SS_{XY}=\sum X_iY_i-\dfrac{1}{n}(\sum X_i)(\sum Y_i)=20485 -\dfrac{247\times 486}{6}=478"


Therefore, based on the above calculations, the regression coefficients (the slope b, and the y-intercept a) are obtained as follows:


"b=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{478}{1240.83}=0.3852\\\ \\a=\bar Y-\bar X\cdot b=81-41.16\times 0.3852=65.14"


Therefore, we find that the regression equation is:


"y=a+bx\\y=65.14+0.3852x"